The real solutions of the inequality log2(x2-1)-log2(x-2)>2 are?


1
Expert's answer
2021-07-19T17:11:39-0400

It follows that "x^2-1>0, x-2>0" and "\log_2(\frac{x^2-1}{x-2})>\log_24." Then "x^2>1, x>2" and "\frac{x^2-1}{x-2}>4." The last inequality is equivalent to "\frac{x^2-1}{x-2}-4>0," which is equivalent to "\frac{x^2-4x+7}{x-2}>0," and hence to "\frac{(x-2)^2+3}{x-2}>0." Since "(x-2)^2+3>0" for any real x"," we conclude that the last inequality is equivalent to the inequality "x>2." Taking into account that the inequality "x^2>1" is equivalent to "|x|>1," and hence "x\in(-\infty,-1)\cup(1,+\infty)," we conclude that the real solutions of the inequality "\log_2(x^2-1)-\log_2(x-2)>2" are all real numbers that are greater than 2.